The ratio of lights to bulbs is 3:5 The number of lights is 24 times the number of bulbs find the number of lights?
1
Expert's answer
2019-11-08T09:06:58-0500

The question is interpreted in the following way.

The ratio of lights to bulbs is 3:5. The number of lights is 24 less than the number of bulbs, find the number of lights?

Solution:

Let

the number of lights be x,

the number of bulbs be y.

According to the task,

the system of equations is

x/y=3/5,

x=y-24.

It follows from the first equation that

x=3y/5

Substituting for x into the second equation

3y/5=y-24

3y/5-y=-24

3y-5y=-120

-2y=-120

y=60

Substituting for y=60 into the second equation

x=60-24

x=36

Answer:

The number of lights is 36.